OpenStudy (goformit100):

Which of the following is not an equivalent form of the compound inequality x + 15 less than or greater to 12 and x + 15 greater than or equal to 5 12 greater than or equal to x + 15 greater than or equal to 5 -3 greater than or equal to x greater than or equal to -10 A number line with a closed circle on -10, a closed circle on -3, and shading in between. A number line with a closed circle on -10, shading to the left, and a closed circle on -3, shading to the right.

OpenStudy (goformit100):

@kelliegirl33

OpenStudy (mathstudent55):

\(x + 15 \le 12\) and \(x + 15 \ge 5\) Subtract 15 from both sides in both inequalities: \(x \le -3\) and \(x \ge -10\)

OpenStudy (anonymous):

x + 15 >= 5 x >= 5 - 15 x >= -10 x + 15 <>12 -- I don't understand the " less then or greater to" x <> 12 - 15 x <> -3 I don't know how to answer this with the " less then or greater to "

OpenStudy (anonymous):

@sasogeek ...can you look at this please

OpenStudy (mathstudent55):

I think he meant "less than or _equal_ to." Notice he used "to" instead of "than."

OpenStudy (anonymous):

well that definitely makes a difference

OpenStudy (anonymous):

x + 15 <= 12 x <= 12 - 15 x <= -3 so we have x <=-3 and x >= -10

OpenStudy (anonymous):

@mathstudent55 ...can you help me with this ? If I go by this I know the last choice is incorrect. But I don't understand where the first two answer choices come in.

OpenStudy (anonymous):

second choice is correct

OpenStudy (goformit100):

ok...

OpenStudy (mathstudent55):

The first choice is: A. \(12 \ge x + 15 \ge 5\) which is the same as \( 5 \le x + 15 \le 12 \)

OpenStudy (mathstudent55):

That is exactly what the original statement means, so choice A is correct.

OpenStudy (anonymous):

b is correct

OpenStudy (anonymous):

c is correct

OpenStudy (goformit100):

And Finally I am confused.

OpenStudy (anonymous):

d is your answer

OpenStudy (anonymous):

if its greater then,shading goes to the right....if it is less then, shading goes to your left

OpenStudy (mathstudent55):

B. is correct because \(-3 \ge x \ge -10 \) means the same as my solution above, \(x \le - \) and \(x \ge -10 \), or \(-10 \le x \le -3\)

OpenStudy (anonymous):

I believe the answer is D...what you think mathstudent ?

OpenStudy (mathstudent55):

Choice C. is correct because it means \(-3 \le x \le -10\)

OpenStudy (mathstudent55):

Choice D. means \(x \le -10 \) or \( x \ge -3 \) which is not the solution.

OpenStudy (anonymous):

thanks for the help mathstudent.....the first choice confused me

OpenStudy (anonymous):

and so did the less then or greater to

OpenStudy (mathstudent55):